Shift report 20011002 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 02 2001 - 23:53:28 EDT

  • Next message: Ian Bearden: "RE: dN/deta, Asymmetry, Background (Re: dndeta figures for impacients )"

    ------------------------------------------------------------
    Date: Tue Oct  2 23:53:27 2001, Shift: 16:00-24:00
    Supervisor: dana beavis9, with: 
    Log book#9 pages 77 - 80
    
    Shift Summary:
    ------------------------------------------------------------
    No usable beam.
    A large beam loss may have caused an increase in pressures
    in the beam pipe.  Next store had high backgounds and did 
    not to run.  Need to check next store.
    May need to stay off for a shift or so.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 02 2001 - 23:54:03 EDT